### v1

# Quantum circuit complexity and unsupervised machine learning of topological order

Abstract
Inspired by the close relationship between Kolmogorov complexity and unsupervised machine learning, we explore quantum circuit complexity, an important concept in quantum computation and quantum information science, as a pivot to understand and to build interpretable and efficient unsupervised machine learning for topological order in quantum many-body systems. To span a bridge from conceptual power to practical applicability, we present two theorems that connect Nielsen’s quantum circuit complexity for the quantum path planning between two arbitrary quantum many-body states with fidelity change and entanglement generation, respectively. Leveraging these connections, fidelity-based and entanglement-based similarity measures or kernels, which are more practical for implementation, are formulated. Using the two proposed kernels, numerical experiments targeting the unsupervised clustering of quantum phases of the bond-alternating XXZ spin chain, the ground state of Kitaev’s toric code and random product states, are conducted, demonstrating their superior performance. Relations with classical shadow tomography and shadow kernel learning are also discussed, where the latter can be naturally derived and understood from our approach. Our results establish connections between key concepts and tools of quantum circuit computation, quantum complexity, and machine learning of topological quantum order.
I Introduction
The discovery of topological phases of matter has opened a new chapter in modern physics, ranging from symmetry-protected band topologies with short-range quantum entanglement such as topological insulators and superconductors Hasan and Kane (2010); Qi and Zhang (2011), to topological quantum orders Zeng and Wen (2015); Zeng et al. (2015); Kitaev and Preskill (2006); Levin and Wen (2006); Kitaev (2003); Semeghini et al. (2021); Satzinger et al. (2021); Google Quantum AI and Collaborators (2024) featured by distributed non-local quantum entanglement such as in Kitaev’s toric code Kitaev (2003); Satzinger et al. (2021). Defying Landau’s paradigmatic approach, where phases and phase transitions are identified through local order parameters (that is, linear functionals of the density matrix of the phase and a local observable), topological phases encode the phase information in global –topological– system properties. This defining difference renders the detection and measurement of topological order challenging for both generic simulation and experimental demonstration.
Recent progress has shown the potential of machine learning for classifying topological phases of matter Carleo and Troyer (2017); Carrasquilla and Melko (2017); Van Nieuwenburg et al. (2017); Zhang and Kim (2017); Liu and van Nieuwenburg (2018); Zhang et al. (2018); Sun et al. (2018); Carleo et al. (2019); Ming et al. (2019); Rem et al. (2019); Lian et al. (2019); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020); Long et al. (2020); Greplova et al. (2020); Balabanov and Granath (2020); Yu and Deng (2021); Mendes-Santos et al. (2021); Leykam and Angelakis (2023); Smolina et al. (2024); Das Sarma et al. (2019). In particular, the supervised approach to quantum many-body physics problems, which can resort to well established theories of learnability and generalizability Valiant (1984); Mohri et al. (2018); Jacot et al. (2018), has resulted in the formulation of rigorous guarantees Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). However, in many practical scenarios, where a priori knowledge of the different phases of matter (and accordingly their labels) does not exist, only unsupervised machine learning can provide access to the desired information. Despite the practical relevance of unsupervised learning, a rigorous theory, tailored to topological quantum order and offering both interpretability and generalizability, has not yet been fully explored.
One prevalent approach to unsupervised learning is based on kernel methods Mohri et al. (2018); Jacot et al. (2018); Rodriguez-Nieva and Scheurer (2019), where feature vectors of the data set are encoded into a kernel in terms of their overlaps or similarities, followed by a nonlinear dimensionality reduction for clustering in a low-dimensional representation space, e.g., the kernel principal component analysis (PCA) Schölkopf et al. (1998); Mika et al. (1999) or the diffusion map algorithm Coifman et al. (2005); Nadler et al. (2006). In the context of topological quantum order, the task then is to formulate a kernel that encodes the nonlinear features of topological quantum states and represents topological similarities or distances by exploiting the primitive notion of topological equivalence. Topologically equivalent quantum states, by definition, can be continuously transformed into each other via a smooth unitary transformation generated by a local Hermitian operator with bounded operator norm Zeng et al. (2015). In this spirit, a kernel based on a path-finding algorithm has been proposed in Scheurer and Slager (2020) and shown good performance in clustering symmetry-protected band topologies, which are characterized by short-range quantum entanglement. Other topological similarity measures, that focus on the closing of the spectral gap at topological transitions, have also been proposed and demonstrated Che et al. (2020); Long and Zhang (2023).
Despite the elegance of these methods in identifying band topologies, the increasing complexity and quantum entanglement in strongly interacting quantum many-body systems may undermine the efficiency and performance of these algorithms. For instance, the path-finding algorithm, when applied to interacting lattice models in order to identify a smooth unitary path generated by a local operator while keeping the parent Hamiltonian gapped, can be principally inconclusive due to the undecidability of the spectral gap for generic quantum many-body Hamiltonians Cubitt et al. (2015).
The inherent problems of previous kernel formulations call for a new kernel ansatz, an ansatz which is able to capture the topological distance and entanglement patterns at different scales while at the same time remaining interpretable and efficiently computable (see Fig. 1 for a schematic of the current state regarding the application of machine learning to the classification of topological phases of matter). To this end, we here adopt the alternative but equivalent perspective that two topologically equivalent multi-qubit states can be transformed into each other via a constant-depth quantum circuit composed of geometrically local quantum gates Zeng et al. (2015); Zeng and Wen (2015), i.e., the smooth unitary transformation that connects them can be decomposed into a shallow quantum circuit. This implies that it is algorithmically trivial or cheap to generate one quantum state given the access to another in the same topological class by fully exploiting the shared structure and entanglement patterns between them.
Figure 1: Landscape of machine learning topological phases of matter. Supervised learning requires prior information or labels, whose efficiency guarantees have been extensively investigated Huang et al. (2022); Lewis et al. (2024); Che et al. (2024); Rouzé et al. (2024); Cho and Kim (2024); Zhao et al. (2024); Onorati et al. (2023); Du et al. (2024); Wanner et al. (2024); Šmíd and Bondesan (2023). Under the constraint of geometric locality Lewis et al. (2024); Rouzé et al. (2024) or with bounded number (No.) of parameters (param.) Che et al. (2024), provably efficient machine learning of quantum many-body systems has been established, e.g., in the framework of the probably approximately correct (PAC)-learnable Valiant (1984); Che et al. (2024). Supervised learning is conceptually more straightforward due to the existence of well-established statistical learning theories. In contrast, unsupervised learning, whose theory is more complicated and less addressed, becomes practically relevant in case of no access to prior labels. The symmetry-protected topological (SPT) order features short-range entanglement, while the topological quantum order exhibits long-range entanglement. The unsupervised learning with the path-finding algorithm Scheurer and Slager (2020) and kernels focusing on closing gaps at topological transitions Che et al. (2020); Long and Zhang (2023), respectively, have shown good performance and interpretability for SPT orders or band topologies. However, the direct generalization of these methods to more complex quantum systems with increased complexity and quantum entanglement faces challenges, due to the intrinsic hardness of identifying (or even the undecidability Cubitt et al. (2015) of) the spectral gap for generic quantum many-body Hamiltonians (See more elaborations and references in the introduction of the main text). Note that the neural network representation of quantum states and reinforcement learning for quantum systems are not included in this landscape.
The correspondence between topological equivalence of quantum states and shallow quantum circuit transformations indicates that, a complexity measure may be related to the topological distance. This is similar to the situation with Kolmogorov complexity, which represents an informational distance between two strings or objects Bennett et al. (2010), and may offer a basis for theoretically optimal solution to unsupervised machine learning Ily (See more details in the following section). We then assert that, the (conditional) generation and prediction of quantum states with the minimal quantum circuit cost may offer a theoretically optimal solution to the unsupervised learning of topological order. The complexity of the generating circuit can be interpreted as a quantum informational distance measure that captures the entanglement change at different scales. Specifically, we propose to use the quantum circuit complexity (QCC) of the unitary path connecting two quantum states as a a suitable distance measure that can be used to construct a kernel for the unsupervised learning of topological order. While the QCC is already a central concept in quantum information theory Chi-Chih Yao (1993); Brandão et al. (2021); Haferkamp et al. (2022); Yi et al. (2024); Liu et al. (2020a), quantum field theory Chapman et al. (2018); Hackl and Myers (2018); Khan et al. (2018) and quantum gravity Susskind (2014); Stanford and Susskind (2014); Brown et al. (2016a, b, 2023), its connection with, and relevance for, the unsupervised machine learning of topological phases of matter has remained mostly unexplored.
In this work, we use the QCC as a pivotal tool to understand and construct unsupervised machine learning of topological order for quantum many-body systems, in an interpretable and efficient manner. Although the exact QCC is generically intractable, we identify various quantities which are upper bounded by the QCC while remaining implementable, and which thus can be efficiciently utilized for classifying topological quantum many-body phases. Specifically, we establish entanglement-based and fidelity-based kernels, respectively. Numerical experiments with a bond-alternating XXZ qubit chain and Kitaev’s toric code model, respectively, demonstrate the excellent performance of these kernels.
II Quantum path planning, topological order and quantum circuit complexity
As mentioned above, the unsupervised machine learning of band topology has been extensively investigated. Band topologies, or the more general symmetry-protected topological (SPT) orders, such as those exhibited by topological insulators, are characterized by the absence of long-range quantum entanglement. This implies that, if the constraint of preserving the symmetry is dropped, SPT quantum states can be transformed smoothly into product states. In contrast, topological quantum order refers to quantum states with long-range quantum entanglement Zeng et al. (2015); for instance, the ground state of Kitaev’s toric code exhibits topological quantum order that manifests in non-zero topological entanglement entropy Kitaev and Preskill (2006); Levin and Wen (2006) and global logical operations Kitaev (2003), while SPT phases are confined to short-range entanglement. Thus, the task of identifying topological quantum order with unsupervised learning presupposes to construct kernel (similarity) functions with feature maps that capture quantum entanglement at a variety of scales. For this purpose, we start by defining the quantum path planning between two multi-qubit quantum states, which lays the ground for our search for a distance measure of clustering topological orders.
Definition 1.
[Quantum path planning (QPP).] The QPP problem is defined as: Given two arbitrary pure $n$-qubit quantum states $\rho_{0}$ and $\rho_{1}$, find continuous paths $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$ $(s\in[0,1])$ that connect $\rho_{0}$ and $\rho_{1}$, where the generator $G(s)$ of $U(s)$ is a hermitian operator for all $0\leq s\leq 1$, and $\|\partial_{s}U(s)\|_{\infty}=\|G(s)\|_{\infty}<+\infty$. In general, we have that $U(s)\in SU\left(2^{n}\right)$ and $G(s)\in su\left(2^{n}\right)$, with the unitary given by
$U(s)$ $=$ ${\cal{P}}\ \mathrm{exp}\left(-i\int_{0}^{s}G(\tau)\mathrm{d}\tau\right),$ (1)
where $U(s=0)=\mathrm{Id}$ is the identity and $\cal{P}$} denotes the path ordering operator.
Note that, for two topologically equivalent gapped quantum states, the generator $G(s)$ in the above QPP must be additionally a geometrically local operator, which underlies why the two quantum states can be transformed into each other via a constant-depth quantum circuit Zeng et al. (2015) (See the following definition).
Definition 2.
(Topological equivalence Zeng and Wen (2015); Zeng et al. (2015).) We say that two gapped ground states $\rho_{0}$ and $\rho_{1}$ of local Hamiltonians $H_{0}$ and $H_{1}$, respectively, are in the same topological phase if there is a local unitary that maps the two states into each other without closing the energy gap. In other words, there is a smooth path of local Hamiltonians $H(s)$ that connects the two parent Hamiltonians with $H(s=0)=H_{0}$ and $H(s=1)=H_{1}$ for $s\in[0,1]$. The Hamiltonian $H(s)$ is gapped above its ground state $\rho(s)$ for all $s\in[0,1]$, where $\rho(s)$ smoothly connects $\rho_{0}$ and $\rho_{1}$, which establishes a smooth path between the two quantum many-body states. In Bachmann et al. (2012), it is show that such a path can be realized with a smoothly parametrized unitary $U(s)$ whose generator $G(s)$ is also a local operator, $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$, and naturally $\rho_{1}=U(s=1)\rho_{0}U^{\dagger}(s=1)$. This is further elaborated in Appendix B.
The above Definition 2 is equivalent to the statement that the two topologically equivalent states can be deformed into each other by a constant-depth quantum circuit Zeng et al. (2015), each layer of which can be divided into geometrically local quantum gate operations. Therefore, the two quantum states differ from each other only in entanglement patterns at short-range scales, while they are equivalent in their long-range entanglement structure.
Intuitively, the depth of the quantum circuit quantifies the scale of the change of the quantum entanglement that is caused by applying the circuit on the reference quantum state $\rho_{0}$. Then, the quantum circuit complexity (QCC) for realizing $U(s=1)$ can serve as a measure of the algorithmic information for generating the target state $\rho_{1}$ from a reference state $\rho_{0}$. As is explained above, this is similar to, and inspired by, the (conditional) Kolmogorov complexity $K(x|y)$, which is the length of the shortest program on a universal computer that outputs a string or dataset $x$ given the access of another string or dataset $y$, and as such represents an algorithmic information distance Bennett et al. (2010) or the shared structure between the two objects Ily . We adopt and adapt this here, using a geometric lower bound of the quantum circuit cost, the Nielsen’s QCC $\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)$, as a distance measure for clustering topological states with distinct entanglement patterns, whose definition is given in the following.
Definition 3.
(Nielsen’s QCC.) The quantum circuit complexity of the unitary $U(s=1)$ can be evaluated through Nielsen’s geometric approach Nielsen (2005); Nielsen et al. (2006). For qubit systems, if we write $G(s)=\sum_{\sigma}h_{\sigma}(s)\ \sigma$, where $\sigma\in P_{n}=\{I,X,Y,Z\}^{\otimes n}$ is a $n$-qubit Pauli operator ($I$ is the $2$-dimensional identity operator and $X,Y,Z$ are the three single-qubit Pauli operators, respectively), and $h_{\sigma}(s)$ denote real coefficients, the Nielsen’s quantum circuit complexity of the first order reads
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\underset{h(s)}{\mathrm{inf}}\ \int^{1}_{0}\sum_{\sigma}\left|h_{\sigma}(s)\right|\ \mathrm{d}s.$ (2)
The original Nielsen’s QCC can be generalized to systems other than qubits Hackl and Myers (2018); Khan et al. (2018); Eisert (2021), where an orthonormal operator basis $\{O_{i}\}$ replaces the $n$-qubit Pauli operators $P_{n}$, with the operator norm $\|O_{i}\|_{\infty}\leq 1$ ($\forall i$). Then the definition of Nielsen’s QCC above remains unchanged.
With the Nielsen’s QCC of geometrically local quantum circuit as a distance measure, the proposed theoretically optimal kernel for the unsupervised learning of topological order is given by
$\mathcal{K}\left(\rho_{0},\rho_{1}\right)=\mathrm{exp}\left[-\beta\ \mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\right],$ (3)
up to a normalization constant, where $\beta>0$ is a hyperparameter. The unitarity of the quantum circuit indicates that Nielsen’s QCC is symmetric under permutations of the two quantum states, i.e., $\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{N}}\left(\rho_{1}\rightarrow\rho_{0}\right)$, which leads to a symmetric kernel in (3).
III Unsupervised machine learning of topological order derived from quantum circuit complexity
While it is in general hard to exactly compute $\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)$ for the QPP, we now establish some approximate substitutes, which both capture essential aspects of Nielsen’s QCC and at the same time are easier to implement in practice. The two theorems that we present in the following provide rigorous guarantees for this purpose.
First, we define the quantum Fisher complexity (QFC) related to the QPP, denoted by $\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)$, which coincides with the Bures distance between the two quantum states given by a fidelity measure and can be efficiently calculated in practice.
Definition 4.
(Quantum Fisher complexity (QFC).) Defining the Uhlmann-Jozsa fidelity between two arbitrary quantum states (pure or mixed) as Nielsen and Chuang (2012)
$F\left(\rho,\tilde{\rho}\right)=\mathrm{tr}\sqrt{\sqrt{\rho}\tilde{\rho}\sqrt{\rho}},$ (4)
where $\mathrm{tr}$ denotes the trace operation, then the corresponding squared Bures distance reads Braunstein and Caves (1994); Liu et al. (2020b); Taddei et al. (2013)
[…]
For a family of smoothly parametrized quantum states $\rho(s)$ by $s\in[0,\,1]$, the quantum Fisher information (QFI) of the quantum state with respect to the parameter $s$ is given by Helstrom (1976); Holevo (1982); Liu et al. (2020b) $\mathcal{F}_{Q}\left(s\right)=\mathrm{tr}\left[\rho\left(s\right)\mathcal{L}^{2}\right]$, with the symmetric logarithmic derivative (SLD) $\mathcal{L}$ determined by $\partial_{s}\rho\left(s\right)=\frac{1}{2}\left[\rho\left(s\right)\mathcal{L}+\mathcal{L}\rho\left(s\right)\right]$. The QFI, which plays a central role in quantum parameter estimation and quantum metrology, can be related to the infinitestmal Bures distance Braunstein and Caves (1994) with
$\frac{1}{4}{\cal{F}}_{Q}(s)\mathrm{d}^{2}s=D^{2}_{B}\left[\rho(s),\rho(s+\mathrm{d}s)\right],$ (6)
representing a susceptibility of the fidelity.
The QFC of the target quantum state $\rho_{1}$ with respect to the reference state $\rho_{0}$ is given by Chapman et al. (2018)
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\underset{G(s)}{\mathrm{inf}}\ \frac{1}{2}\int^{1}_{0}\sqrt{{\mathcal{F}}_{Q}(s)}\ \ \mathrm{d}s,$ (7)
which represents the optimal Bures distance between the two quantum states Taddei et al. (2013), reflecting the overall fidelity variation, as indicated by (6). Note that we have a symmetric QFC under permmutations of its inputs, i.e., $\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)=\mathcal{C}_{\mathcal{F}}\left(\rho_{1}\rightarrow\rho_{0}\right)$.
With the QFC and the Bures distance, we are ready to formulate the following theorem.
Theorem 1.
(Quantum circuit complexity upper bounds fidelity change.)  The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ is lower bounded by the QFC and the Bures distance,
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)$ $\geq$ ${\cal{C}}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)$ (8)
$\geq$ $\frac{D_{B}\left(\rho_{0},\rho_{1}\right)}{\sqrt{2}}.$
Furthermore, the QCC of a geometrically local quantum circuit for the same QPP is approximately lower bounded by
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\gtrsim\frac{1}{\sqrt{2}}\,\underset{\Delta}{\sum}\,D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right],$ (9)
where $\rho\left(\Delta\right)$ is the reduced density matrix supported on the subsystem $\Delta$ of constant size, and the summation goes over non-overlapping subsystems which, together with their neighboring environments, cover the whole system.
The proof of Theorem 1 is given in Appendix A. Note that, without the constraint of geometric locality of the quantum circuit, the overall Bures distance in (8) is equivalent to the $L_{2}$-norm distance between the two density matrices, which has been shown not to be a suitable distance measure for topological phases Che et al. (2020).
Under the geometric locality of the quantum circuit, we exploit Theorem 1 to construct a fidelity-based kernel. The Uhlmann-Jozsa fidelity between the $r$-body reduced density matrices with $r=\{1,2,\cdots,R\}$ can be used for this purpose, where $R$ is a scale of cutoff. For instance, for qubits hosted on a lattice $\Lambda=[1,\ L]^{D}$ of lattice size $L$ and dimension $D$, the proposed kernel based on fidelity reads (up to some normalization constant)
$\mathcal{K}_{\mathrm{F}}\left(\rho,\tilde{\rho}\right)=\mathrm{exp}\left\{\beta\ \underset{r=1}{\sum^{R}}\omega_{r}\underset{\Delta\in P_{r}(\Lambda)}{\sum}F\left[\rho(\Delta),\ \tilde{\rho}(\Delta)\right]\right\},$ (10)
where $P_{r}(\Lambda)=\{\Delta|\Delta\subset\Lambda\ \text{and}\ \left|\Delta\right|=r\}$, with $\left|\Delta\right|$ being the size (cardinality) of the sublattice $\Delta$; $\rho(\Delta)=\mathrm{tr}_{(\Lambda-\Delta)}\rho$ is the reduced density matrix supported on $\Delta$; $\omega_{r}$ is a weight function; and $F(\cdot,\cdot)$ is the Uhlmann-Jozsa fidelity. While the dimension of the reduced density matrix depends exponentially on its size $r$, this fidelity-based kernel can be efficiently estimated in practical implementations for $R=\mathrm{const.}$ (For instance, a small $r=2$ is already sufficient in many relevant applications. See Theorem 1 and the section on numerical experiments below.)
The second approximation of the Nielsen’s QCC that we propose focuses on the entanglement change. Denoting by $\rho$ a pure density matrix of a $n$-qubit chain (rearranged from an arbitrary spatial geometry following a specified order), and $A_{k}=\{1,2,...,k\}\subset[n]$ a subset of the system, which is obtained from a cut at $k\in[1,\ n-1]$, then the reduced density matrix supported on $A_{k}$ reads $\rho(A_{k})=\mathrm{tr}_{B_{k}}\rho$, where $B_{k}=\{k+1,k+2,...,n\}$. With the entanglement entropy over the cut $k$ given by
$S_{k}(\rho)=\mathrm{tr}\left[\rho(A_{k})\ln\rho(A_{k})\right],$ (11)
we have the following theorem:
Theorem 2.
[Quantum circuit complexity upper bounds entanglement change (A generalized version of Observation 1 in Eisert (2021)).] The QCC for the QPP from $\rho_{0}$ to $\rho_{1}$ with a geometrically local quantum circuit (with respect to the $n$-qubit chain) is lower bounded by
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\geq\frac{c}{\left(n-1\right)}\sum^{n-1}_{k=1}\left|S_{k}(\rho_{1})-S_{k}(\rho_{0})\right|,$ (12)
for some constant $c>0$.
The proof of Theorem 2 is given in Appendix A. Theorem 2 can be readily generalized to situations where the qubits are arranged in higher spatial dimensions and other geometries, provided that the quantum circuit is applied in a geometrically local manner accordingly. Nielsen’s QCC is then lower bounded by the entanglement change distributed over suitable cuts covering non-trivial bonds of the system.
Theorem 2 inspires an entanglement-based kernel, given by
[…]
which is particularly useful when the entanglement profile over nontrivial bonds can be efficiently estimated. This is, for instance, the case in tensor-network simulations of quantum many-body states such as the density matrix renormalization group (DMRG) White (1992); Schollwoeck (2011), or in experimentally prepared quantum states measured via classical shadow tomography Huang et al. (2020).
Both lower bounds of Nielsen’s QCC in (9) and (12) potentially capture important information about the topological order. The fidelity-based kernel with reduced density matrices of constant size can be efficiently calculated and benchmarked (For instance, with classical shadow representations, see the following section). The entanglement-based kernel is more stringent and contains more information about quantum entanglement at multiple scales.
IV Relation to classical shadows
Classical shadow (CS) tomography is an experimentally friendly approach to access many properties of quantum many-body states with few-shot measurements Huang et al. (2020). In this formalism, the quantum state is approximated by its CS estimator,
$\rho\approx S_{T}\left(\rho\right)=\frac{1}{T}\sum_{t=1}^{T}\sigma^{(t)}_{1}\otimes\sigma^{(t)}_{2}\cdots\otimes\sigma^{(t)}_{n}$, (14)
in a total of $T$ measurements, where $\sigma^{(t)}_{i}$ is the local CS representation under the Pauli measurement for the $i$-th qubit in the $t$-th measurement. The shadow kernel suggested in Huang et al. (2022), given by
[…]
can be understood as a special case of the fidelity-based kernel in (10), if the quantum state $\rho$ and the $r$-body reduced density matrices are represented by their classical shadows, if the weight function is taken to be $\omega_{r}=\nu^{r}/r!$, and if we take the limit $R\rightarrow\infty$, where $\beta>0$ and $\nu>0$ are constant hyperparameters. In this case, density matrices with small values of $r$ dominate in the summation, and randomized measurements reduce the complexity in calculating fidelity-based kernels.
[…]
V Numerical experiments
In this section, we provide numerical demonstrations for the developed theory of unsupervised machine learning of topological order based on the QCC.
V.1 The bond-alternating XXZ spin chain
We begin with a benchmark model, the bond-alternating XXZ spin-$\frac{1}{2}$ chain Pollmann and Turner (2012); Elben et al. (2020); Huang et al. (2022), whose Hamiltonian is given by
$H_{\mathrm{XXZ}}=\underset{i=1}{\sum^{n-1}}J_{i}\left(X_{i}X_{i+1}+Y_{i}Y_{i+1}+\delta Z_{i}Z_{i+1}\right)-h_{0}\underset{i=1}{\sum^{n}}Z_{i},$ (16)
where $J_{i}=J_{1}$ if $i$ is odd and $J_{i}=J_{2}$ if $i$ is even, respectively; $\delta$ is a detuning parameter and $h_{0}$ denotes the strength of the external field.
In Fig. 2, we present a comparison between the fidelity- and entanglement-based kernels for this model, respectively, together with the nonlinear embedding into a two-dimensional representation space from the diffusion map algorithm Coifman et al. (2005); Nadler et al. (2006); Rodriguez-Nieva and Scheurer (2019); Che et al. (2020); Scheurer and Slager (2020). The dataset consists of ground-state properties (fidelities and entanglements) obtained from DMRG calculations of the model. Fixing $h_{0}=0$, $\delta=3$ and varying the value of $J_{2}/J_{1}$, we find that the unsupervised learning produces distinct clusters that can be associated with three phases (trivial, symmetry broken, and topological), in good agreement with independent calculations of topological invariants Pollmann and Turner (2012); Elben et al. (2020); Huang et al. (2022). More details can be found in the caption of Fig. 2. While we are ultimately aiming at the unsupervised learning of topological order in the presence of non-local quantum entanglement, Fig. 2 shows that the fidelity- and entanglement-based kernels also perform well when clustering symmetry-broken phases and symmetry-protected topological orders Elben et al. (2020) that are characterized by short-range entanglement.
[…]
V.2 Kitaev’s toric code
As a second demonstration, we cluster the ground state of Kitaev’s toric code with random product states, where the former has topological quantum order while the latter is topologically trivial. The toric code Hamiltonian defined on a $\mathrm{2D}$ lattice of size $L_{x}\times L_{y}$ is given by Kitaev (2003)
$H_{\mathrm{TC}}=-\underset{v}{\sum}A_{v}-\underset{p}{\sum}B_{p},$ (17)
where $v$ denotes a vertex and $p$ is the face of the lattice, respectively; $A_{v}=\underset{j\in star(v)}{\prod}X_{j}$ and $B_{p}=\underset{j\in boundary(p)}{\prod}Z_{j}$ are the Kitaev’s star and face operators, respectively. The number of qubits for this model is $n=2L_{x}L_{y}$. The toric code ground state is gapped and exhibits area-law entanglement Kitaev and Preskill (2006). Therefore, for the purpose of a proof-of-principle demonstration and in order to obtain the entanglement profile for the kernel with a small system size, we use a MPS representation and DMRG to find a reasonably accurate solution for the ground state. For instance, numerical experiments show that, for a $4\times 4$ lattice with periodic boundary conditions in both dimensions, a maximal MPS bond dimension of $64$ is sufficient to produce accurate results.
[…]
In Fig. 3 we plot a comparison between the unsupervised machine learning of the toric code ground state (blue dots) and the random product states (RPS) (red dots) of qubits (i.e., random bit strings of spin ups and downs), with fidelity- and entanglement-based kernels, respectively. Both kernels perform well in clustering the topologically quantum ordered state and the trivial phase, while the entanglement-based one is more robust under application of two-qubit random unitaries (with respect to the Haar measure) to both the toric code and the RPS. Figure 3 (a) is similar to the result obtained from the shadow kernel learning in Huang et al. (2022), confirming the relation to classical shadows that was elaborated above. The entanglement created by applying the two-qubit random circuit grows exponentially with the circuit depth. More details about the plot can be found in the caption of Fig. 3.
V.3 The extended toric code
The original toric code is defined on a torus with periodic boundary conditions in both spatial dimensions. In contrast, the extended toric code (ETC) is defined on an infinite cylinder (e.g., the cylinder axis is along the $x$ direction and the $y$ direction is periodic with circumference $L_{y}$), and additionally features the Wilson and t’Hooft loop operators $W$ and $H$, respectively. The corresponding Hamiltonian is given by Hauschild and Pollmann (2018)
$H_{\mathrm{ETC}}=H_{\mathrm{TC}}-J_{\mathrm{W}}W-J_{\mathrm{H}}H-h\underset{i=1}{\sum^{n}}Z_{i},$ (18)
where $H_{\mathrm{TC}}$ is the original toric code Hamiltonian in (17), and $h$ is a (uniform) local field strength. The loop operators are defined as products
$W=\underset{i\in E_{v}}{\prod}Z_{i},\,\,\,\,\,H=\underset{i\in E_{h}}{\prod}X_{i}$ (19)
around the cylinder through vertical edges ($E_{v}$) and horizontal edges ($E_{h}$) of the lattice graph, respectively.
In Fig. 4, by fixing the values of $J_{W}=+1$ and $J_{H}=-1$, which stabilizes the ground state into one of the four degenerate sectors, and by varying the strength of the local perturbation field $h$, we plot a comparison between the unsupervised machine learning of the ETC ground state (blue dots) and the random product states (RPS) (red dots) of qubits (i.e., random bit strings of spin ups and downs), with fidelity- and entanglement-based kernels, respectively. Note that, for finite values of $h$, the ground states are no longer exactly solvable. The required maximal MPS bond dimension increases rapidly with larger values of $h$ even for small system sizes (For the system size and the parameter range we adopted in Fig. 4, a maximal bond dimension of $600$ is sufficient to produce accurate results). The topological clustering shows its stability against local perturbations. More details about the plot can be found in the caption of Fig. 4.
The DMRG ground-state search in the numerical experiments was performed using the TeNPy Library Hauschild and Pollmann (2018), and some example codes there are reused for preparing the dataset of our machine learning model.
VI Discussion and outlook
In summary, we used quantum circuit complexity, a vital concept in quantum computation, to understand and to build viable distance measures and kernels for the unsupervised machine learning of topological order, targeting interpretability and generalizability. By resorting to Nielsen’s quantum circuit complexity, two easier-to-implement kernels based on fidelity and entanglement were proposed and their effectiveness numerically verified.
The shadow kernel learning of Huang et al. (2022) results as a natural derivative of our method based on the Nielsen’s QCC, when the fidelity-based similarity and classical shadow representation of the quantum state are used. The alternative kernel based on the entanglement entropy profile encodes more information about entanglement patterns in topological orders, and thus exhibits superior robustness against two-qubit random noises; moreover, they are more interpretable in terms of the relationship between topological quantum orders and long-range entanglement. Recent theoretical studies Vermersch et al. (2024a, b) have proposed protocols for the efficient measurement of many-body entropies, which indicates that the entanglement-based kernel can be calculated efficiently from experimental data. This implies that our method can be readily applied to experimentally prepared multi-qubit quantum states; for instance, data obtained from near-term quantum computing hardware. When limited to SPT order and compared to the recent results in Wu et al. (2024) about neural-network learning topological orders of the bound-alternating XXZ model, our method is fully unsupervised and results in sharper phase boundaries in clustering quantum phases of this model while providing good interpretability.
In future work, the kernels proposed here may be also used for supervised learning tasks, such as classification or regression of topological phases, with an $\mathcal{O}(1)$ sample complexity. The nonlinear feature vectors (e.g., random Fourier features) given by the entanglement kernel can be used to learn about properties related to entanglement structures of the quantum many-body state. Also, in the demonstrated examples, so far we focused on pure and gapped ground states. Possible extensions could be applications in gappless, topological quantum criticalities, e.g., with the multiscale entanglement renormalization ansatz (MERA) Vidal (2008), or in the unsupervised machine learning of entanglement transitions Ippoliti et al. (2021); Granet et al. (2023), as well as mixed-state topological order Fan et al. (2024); Ellison and Cheng (2025); Sohal and Prem (2025); Wang et al. (2025); Ma and Turzillo (2025).
Our current approach is primarily based on the classical machine learning of topological quantum order by exploiting Nielsen’s QCC as an informational distance. It might also be interesting to explore quantum machine learning Biamonte et al. (2017); Bartkiewicz et al. (2020) in terms of parametrized quantum circuits (e.g.,  Ivashkov et al. (2024)), which may find approximately the minimal quantum circuit matching the reference and target quantum states in the data set. The combination with reinforcement learning for quantum state generation Barr et al. (2020); Yao et al. (2021); Zeng et al. (2023) can be also promising. We also note a very recent work Lake et al. (2025) which uses exact quantum algorithms for topological quantum phase recognition, while restricted to one-dimensional SPT phases.
We believe that the results presented here make, both conceptually and technically, an important step toward an interpretable and generalizable theory of unsupervised machine learning of topological order, and as such will boost the interplay between the research fields of (interpretable) machine learning phases of matter, quantum complexity, quantum parameter estimation, and quantum computation.
VII Acknowledgments
C.G. is partially supported by RIKEN Incentive Research Projects. X.W. is supported by the Hangzhou Joint Fund of the Zhejiang Provincial Natural Science Foundation of China under Grant No. LHZSD24A050001, and the Science Foundation of Zhejiang Sci-Tech University (Grants No. 23062088-Y, and No. 23062181-Y), and NSFC (No. 24062271-A). F.N. is supported in part by: the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2, the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).
Appendix A Proof of Theorems
Proof of Theorem 1: For a pure quantum state $\rho(s)=|\psi(s)\rangle\langle\psi(s)|$ in the unitary evolution defined in the QPP of the main text, the QFI for the parameter $s$ reads
${\cal{F}}_{Q}(s)$ $=$ $4\left(\langle\partial_{s}\psi(s)|\partial_{s}\psi(s)\rangle-\left|\langle\partial_{s}\psi(s)|\psi(s)\rangle\right|^{2}\right)$ (20)
$=$ $4\left\{\mathrm{tr}\left(G^{2}(s)\rho(s)\right)-\left[\mathrm{tr}\left(G(s)\rho(s)\right)\right]^{2}\right\}$
$=$ $4\text{Var}\left[G(s)\right],$
where we have used the relation $|\psi(s)\rangle=U(s)|\psi(0)\rangle$ and the hermitian generator $G(s)=i\left[\partial_{s}U(s)\right]U^{\dagger}(s)$; and Var denotes the variance with respect to the evolved quantum state $\rho(s)$. The variance of the generator $G(s)$ is upper bounded by
$\text{Var}\left[G(s)\right]\leq\frac{1}{4}\left[\lambda_{\text{max}}(s)-\lambda_{\text{min}}(s)\right]^{2}\leq\|G(s)\|^{2}_{\infty},$ (21)
where $\lambda(s)$ is the eigenvalue of $G(s)$; and to obtain the last inequality, we have used $|a-b|\leq|a|+|b|$ for real numbers $a$ and $b$, and that, by definition, the operator norm of the generator $\|G(s)\|_{\infty}=|\lambda(s)|_{\text{max}}$. The first inequality in (21) is saturated when $|\psi(s)\rangle=\frac{1}{\sqrt{2}}\left(|\lambda_{\text{max}}(s)\rangle+|\lambda_{\text{min}}(s)\rangle\right)$, i.e., the quantum state is an equal weight superposition of eigenstates with maximal and minimum eigenvalues of $G(s)$, respectively; and the last inequality is saturated when $\lambda_{\text{max}}(s)=-\lambda_{\text{min}}(s)>0$.
On the other hand, the operator norm of the $G(s)=\sum_{\sigma}h_{\sigma}(s)\ \sigma$ of Nielsen’s circuit complexity is upper bounded by
$\|G(s)\|_{\infty}\leq\sum_{\sigma}|h_{\sigma}(s)|,$ (22)
where we have used the subadditivity of the operator norm and $\|\sigma\|_{\infty}=1$ for $n$-qubit Pauli operators $\sigma$. Therefore, we have
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{\sigma}|h_{\sigma}(s)|$ (23)
for all possible paths $h_{\sigma}(s)$ ($s\in[0,1]$). Then by integrating over $s\in[0,1]$ on both sides of (23) followed by taking the infimum over all possible paths, the first inequality in (8) holds naturally, i.e.,
$\mathcal{C}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right)\leq{\cal{C}}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right).$ (24)
The second inequality in (8),
$\frac{D_{B}\left(\rho_{0},\rho_{1}\right)}{\sqrt{2}}$ $\leq$ ${\cal{C}}_{\mathcal{F}}\left(\rho_{0}\rightarrow\rho_{1}\right),$ (25)
which lower bounds the QFC by the Bures distance between $\rho_{0}$ and $\rho_{1}$, can be readily and easily derived from the result in Taddei et al. (2013).
The above proof can be directly generalized to systems other than qubits. Suppose that the generator of the unitary quantum path planning is
$G(s)=\sum_{i}h_{i}(s)\ O_{i},$ (26)
where $O_{i}$ is an orthonormal operator basis with operator norm $\|O_{i}\|_{\infty}\leq 1$, then, by replacing the summation in the right-hand side of (23) to go over all orthonormal operator bases of the generator, (23) also holds, i.e.,
$\sqrt{\mathcal{F}_{Q}(s)}\leq 2\sum_{i}|h_{i}(s)|.$ (27)
Next, a suitable similarity measure for topological quantum order imposes the additional constraint of geometric locality of the QPP, which requires that $O_{i}$ is a geometrically local operator for all possible values of $i$, i.e., $O_{i}$ is supported on a constant number of neighboring particles. We consider a reduced density matrix $\rho\left(s|\Delta\right)$ supported on a subsystem $\Delta$ of constant size, undergoing a nonunitary evolution along the unitary QPP of the total system generated by $G(s)$. While the relation between the Bures distance and the QFI in (6) still holds for mixed quantum states with the Uhlmann-Jozsa fidelity, the expression of the QFI for $\rho\left(s|\Delta\right)$ with respect to $s$ is more complicated than (20). Now the QFI has the general form given by the symmetric logarithmic derivative (SLD) $\mathcal{L}$, with
$\mathcal{F}_{Q}\left(s|\Delta\right)=\mathrm{tr}\left[\rho\left(s|\Delta\right)\mathcal{L}^{2}\right],$ (28)
where $\mathcal{L}$ is determined by
$\partial_{s}\rho\left(s|\Delta\right)=\frac{1}{2}\left[\rho\left(s|\Delta\right)\mathcal{L}+\mathcal{L}\rho\left(s|\Delta\right)\right].$ (29)
We make a Trotter decomposition of the unitary evolution generated by $G(s)$ into a product of many sufficiently small, finite time steps. At each Trotter time step, we can estimate an upper bound for $\mathcal{F}_{Q}\left(s|\Delta\right)$ of the subsystem. The evolution of the reduced density matrix of the subsystem $\Delta$ results from a partial-trace operation (with respect to the rest of the system) on the total system which undergoes the unitary evolution generated by $G(s)$. The QFI of the reduced density matrix with respect to $s$ does not exceed that of its enlarged pure system, i.e.,
$\mathcal{F}_{Q}\left(s|\Delta\right)\leq\mathcal{F}_{Q}\left(s\right),$ (30)
where the latter can be estimated via (27) through unitary evolution. Note that (30) can be readily obtained with the monotonicity of the Uhlmann-Jozsa fidelity of two quantum states $\rho$ and $\tilde{\rho}$ Nielsen and Chuang (2012), i.e.,
[…]
under the trace-preserving quantum operation $\mathcal{E}$, which means that the Bures distance and QFI are contractive under the partial-trace operation.
Due to the geometric locality of the generator and of the quantum circuit applied to the system as well as the very small evolved time step, the generators $\{O_{i}\}$ that govern the time evolution of the subsystem $\Delta$ are restricted to include only those supported on $\Delta$ and its neighboring environment, where the latter is directly coupled to $\Delta$ by applying the quantum circuit. We denote $\tilde{\Delta}$ as the subsystem $\Delta$ and its neighboring environment, where the size $|\tilde{\Delta}|\leq\text{constant}$. Consequently, the summation on the right-hand side of (27) for upper bounding $\mathcal{F}_{Q}\left(s|\Delta\right)$ must be restricted to the basis operators which are locally supported on $\Delta$ and its neighboring environment, i.e.,
$\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\leq\frac{1}{2}\sqrt{\mathcal{F}_{Q}\left(s\right)}\leq\underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|,$ (32)
where $\mathcal{S}_{i}$ is the support (on the system) of the $i$-th orthonormal basis operator $O_{i}$ in the expansion of the generator $G(s)$.
Then, by integrating over $s\in[0,\,1]$, followed by summing over different such subsystems of constant size to cover the whole system in order to cover the support of the generator, we have
$\underset{\Delta}{\sum}\ \frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s\leq\int_{0}^{1}\ \underset{\Delta}{\sum}\ \ \underset{i\ :\ \mathcal{S}_{i}\subset\tilde{\Delta}}{\sum}|h_{i}(s)|\ \mathrm{d}s\approx\int_{0}^{1}\underset{i}{\sum}|h_{i}(s)|\ \mathrm{d}s,$ (33)
where the summation in the right-hand side of the last approximate equation goes over all operator bases of the generator. Moreover, from the result presented in Taddei et al. (2013), we can derive that
$\frac{1}{\sqrt{2}}D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right]\leq\frac{1}{2}\int_{0}^{1}\sqrt{\mathcal{F}_{Q}\left(s|\Delta\right)}\ \mathrm{d}s,$ (34)
which always holds for arbitrary smooth paths between the two reduced density matrices, caused by the smooth unitary path of the total system followed by the partial-trace operation, i.e., it is path-independent. Then we can take the infimum over all possible paths of the unitary QPP on the right-hand side of the last equation in (33) to obtain a tighter bound, and conclude that Nielsen’s QCC of geometrically local quantum circuits is approximately lower bounded by the summation over fidelity distances of the reduced density matrices supported on non-overlapping subsystems $\{\Delta\}$ which, together with their neighboring environments, cover the total system, i.e.,
$\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right)\gtrsim\frac{1}{\sqrt{2}}\,\underset{\Delta}{\sum}\,D_{B}\left[\rho_{0}\left(\Delta\right),\rho_{1}\left(\Delta\right)\right].$ (35)
(This completes the proof of Theorem 1.)
Proof of Theorem  2: Here we follow a similar procedure as in Eisert (2021), but with more details in order to generalize the result to two arbitrary multi-qubit states in a generic QPP problem. Arrange the $n$-qubit system into a qubit chain $\{1,2,\cdots,n\}$ following a specific order, and take a subsystem over the $k$-th cut, $A_{k}=\{1,2,\cdots,k\}$, and its environment $B_{k}=\{k+1,k+2,\cdots,n\}$ as in the statement before Theorem 2. Further, we assume that the coupling in the generator $G(s)$ of $U(s)$ between $A_{k}$ and $B_{k}$ is spatially local with respect to the qubit chain. By Trotter decomposition, slicing $s\in[0,\,1]$ into small time steps $\{s_{j}\}_{j=0}^{T_{s}-1}$, and at each time interval $\Delta s_{j}=\left|s_{j}-s_{j-1}\right|$ ($j\in[1,T_{s}-1]$), the unitary evolution operator of the QPP is given by
$U(s_{j})=\mathrm{exp}\left[-i\Delta s_{j}G(s_{j})\right].$ (36)
The generators $G(s_{j})$ can be decomposed as
$G(s_{j})=G_{A_{k}}(s_{j})+G_{B_{k}}(s_{j})+G_{k}(s_{j}),$ (37)
where $G_{A_{k}}(s_{j})$ and $G_{B_{k}}(s_{j})$ are supported on $A_{k}$ and $B_{k}$, respectively, and
$G_{k}(s_{j})=\underset{O^{(k)}}{\sum}h_{O^{(k)}}\left(s_{j}\right)O^{(k)},$ (38)
is the local coupling between the subsystems $A_{k}$ and $B_{k}$, with $O^{(k)}$ a geometrically local basis operator at the $k$-th bond of operator norm $\left\|O^{(k)}\right\|_{\infty}\leq 1$. With rigorous results already proved in Bravyi et al. (2006); Bravyi (2007); Hutter and Wehner (2012); Van Acoleyen et al. (2013); Mariën et al. (2016) concerning the Small Incremental Entangling (SIE), the rate of the entanglement over the cut $k$ generated or destroyed in this small time slice is upper bounded by
$\left|\frac{\mathrm{d}}{\mathrm{ds}}S_{k}\left[\rho(s)\right]\right|\,\Bigg{|}_{s=s_{j}}$ $\leq$ $c^{\prime}\,\ln d_{k}\,\left\|G_{k}(s_{j})\right\|_{\infty}$ (39)
$\leq$ $c^{\prime}\,\ln d_{k}\,\underset{O^{(k)}}{\sum}\left|h_{O^{(k)}}\left(s_{j}\right)\right|,$
where $c^{\prime}>0$ is a constant and $d_{k}=\min\{d_{A_{k}},d_{B_{k}}\}\leq 2^{n-1}$ is the minimal Hilbert-space dimension of the subsystems. With (39), integrating over $s\in[0,\,1]$ (summing over all infinitestmal time slices $\Delta s_{j}$) and using the relation $\left|\int_{0}^{1}f(s)\mathrm{d}s\right|\leq\int_{0}^{1}\left|f(s)\right|\mathrm{d}s$ for smooth functions, followed by summing over all cuts $k$, we obtain that Nielsen’s circuit complexity measure is, for all possible paths, lower bounded by the entanglement change averaged over the bonds. We can then take the infimum of the circuit complexity measure over all paths allowed to obtain a tighter bound, i.e.,
$\frac{c}{\left(n-1\right)}\sum^{n-1}_{k=1}\left|S_{k}(\rho_{1})-S_{k}(\rho_{0})\right|\leq\mathcal{C}_{\mathcal{N}}\left(\rho_{0}\rightarrow\rho_{1}\right),$ (40)
for some constant $c=1/\left(c^{\prime}\ln 2\right)>0$. Note that here the reference quantum state $\rho_{0}$ is not restricted to a product state, and we find an averaging prefactor 1/(n1)1/(n-1) compared to the result in Eisert (2021). (This completes the proof of Theorem 2.)
Appendix B Topological equivalence of gapped ground states and quantum circuit complexity of adiabatic quantum computation
As mentioned in Definition 2, two gapped quantum states $\rho_{0}$ and $\rho_{1}$ in the same topological phase can be connected via a smooth path $\rho(s)=U(s)\rho_{0}U^{\dagger}(s)$ for $s\in[0,1]$. In case of gapped spectra, an explicit expression for the generator $G(s)$ is provided in Bachmann et al. (2012), within the spectral flow formalism:
$G(s)$ $=$ $\int_{-\infty}^{+\infty}\mathrm{d}t\ W_{\gamma}(t)\ \mathrm{e}^{itH(s)}\,\partial_{s}H(s)\,\mathrm{e}^{-itH(s)},$ (41)
where $H(s)$ is the Hamiltonian path with a spectral gap $\delta(s)\geq\gamma>0\left(\forall s\in[0,1]\right)$ and with the ground state $\rho(s)$; $W_{\gamma}(t)\in L^{1}(\mathbb{R})$ is a decaying weight function of the spectral flow satisfying $\left\|W_{\gamma}\right\|_{1}\leq c_{0}/\gamma$, for some positive constant $c_{0}$. In this situation, the Nielsen QCC may be explicitly obtained by optimization based on differential geometry, which deserves further investigation.
From the Proof of Theorem 1, we know that the integrand of Nielsen’s QCC is lower bounded by the operator norm of the generator, with $\|G(s)\|_{\infty}\leq\sum_{\sigma}|h_{\sigma}(s)|$. This inequality is already sufficiently satisfied if we replace $\|G(s)\|_{\infty}$ on the left-hand side by its maximal value. With (41) and the properties of the decaying weight function of the spectral flow, one finds the maximal value to satisfy
$\|G(s)\|_{\infty}\leq\frac{c_{0}\|\partial_{s}H(s)\|_{\infty}}{\gamma}.$ (42)
It follows that a lower bound for Nielsen’s QCC can be given by an optimization over $G(s)$ after integrating the right-hand side of (42) over $s\in[0,1]$, which is similar (but not identical) to the adiabaticity condition in adiabatic quantum computation Albash and Lidar (2018).